OpenStudy (anonymous):

What are the solutions of 8x2 = 24x? x = 8, x = 3 x = 8, x = −3 x = 0, x = −3 x = 0, x = 3

OpenStudy (johnweldon1993):

Do you know how to start it?

OpenStudy (anonymous):

I think its D, I believe that would make it an even ration

OpenStudy (johnweldon1993):

You would be correct

OpenStudy (johnweldon1993):

but you know how to make sure it correct right?

OpenStudy (anonymous):

No i'm not sure exactly

OpenStudy (johnweldon1993):

Alright.....so....we would take our original equation 8x² = 24x and move everything to the left side...so we'll have 8x² - 24x = 0 now we would factor the left side....any idea how to do that?

OpenStudy (anonymous):

Honestly dude, I have no idea, I have no clue how I have an A in this class

OpenStudy (johnweldon1993):

haha....well it's easy 8x² - 24x = 0 see how both 8...and 24 are divisable by 8? this means we can factor out an 8 it will look like this 8(x² - 3x) = 0 now look and see that there is also at least 1 'x' in each term in those parenthesis...we can factor out an 'x' as well now it will look like this 8x(x - 3) = 0 so...how do we make that equation = 0? well....either 8x...has to equal 0...meaning x will = 0 or x - 3 has to equal 0....meaning x will equal 3 so the 2 solutions are 0, and 3 or as you have already seen....answer choice D

OpenStudy (anonymous):

@johnweldon1993 You're awesome!! Thank you so much for all of your help!! I wish I could give you more than one medal

OpenStudy (johnweldon1993):

don't worry about it man! glad to help!